If one form of the correct answer of a sum of two rational expressions is (-9)/ (k-9), what would be an alternate form of the answer if the denominator is 9-k?
Answer #1

(-9)/(k-9) can be written as

(-9)/(-(9-k)) i.e., taking negative(-) as common in denominator

The negative in numerator and the negative in denominator will get cancelled. Thus we will get (9)/(9-k)

(-9)/(k-9) = (-9)/(-(9-k)) = (9)/(9-k)

Therefore the alternate form of (-9)/(k-9) is (9)/(9-k)